# suppReferences

### v3

@@ Abstract
Real-world misinformation can be partially correct and even factual but misleading. It undermines public trust in science and democracy, particularly on social media, where it can spread rapidly. High-quality and timely correction of misinformation that identifies and explains its (in)accuracies has been shown to effectively reduce false beliefs. Despite the wide acceptance of manual correction, it is difficult to be timely and scalable, a concern as technologies like large language models (LLMs) make misinformation easier to produce. LLMs also have versatile capabilities that could accelerate misinformation correction—however, they struggle due to a lack of recent information, a tendency to produce false content, and limitations in addressing multimodal information. We propose Muse, an LLM augmented with access to and credibility evaluation of up-to-date information. By retrieving evidence as refutation or supporting context, Muse identifies and explains (in)accuracies in a piece of content—not presupposed to be misinformation—with references. It also describes images and conducts multimodal searches to verify and correct multimodal content. Fact-checking experts evaluate responses to social media content that are not presupposed to be (non-)misinformation but broadly include incorrect, partially correct, and correct posts that may or may not be misleading. We propose and evaluate 13 dimensions of misinformation correction quality, ranging from the accuracy of identifications and factuality of explanations to the relevance and credibility of references. The results demonstrate Muse’s ability to promptly write high-quality responses to potential misinformation on social media—overall, Muse outperforms GPT-4 by 37% and even high-quality responses from laypeople by 29%. This work reveals LLMs’ potential to help combat real-world misinformation effectively and efficiently.
@@ Introduction
Misinformation, which broadly includes partially incorrect and factual but misleading content[1, 2], has far-reaching and detrimental effects on individuals and society[3, 1, 2, 4, 5]. It erodes public trust in government, decreases civil engagement in elections, and has been viewed as a threat to democracy[6, 2, 7]. Evidence has demonstrated that election misinformation helped fuel the January 6th attack on the U.S. Capitol, where five people died and more than 100 police officers were injured[8]. Misinformation also drastically increases during outbreaks and disasters, as seen with the “infodemic” during the COVID-19 pandemic, which significantly increased vaccine hesitancy[9, 10]. The result is tragic, where COVID-19 vaccines could have prevented at least 232,000 of COVID-19-associated deaths between May 2021 and September 2022 in the U.S. alone[11]. Concerns over misinformation on social media have been particularly significant[12, 13, 14], as the social media context interferes with truth discernment, where users post content without professional moderation and often consume news in a hasty and distracted way[15, 16].
@@
Fortunately, high-quality and timely correction of misinformation, which identifies what part(s) of its content is or is not accurate and explains why that part of the content is (in)accurate with references, has been shown to effectively reduce the spread of misinformation and false beliefs[17, 18, 4, 19]. While domain experts (e.g., on FactCheck.org) and groups of laypeople (e.g., on X Community Notes, formerly Twitter Birdwatch) have played pivotal roles in correcting misinformation[19, 20, 17], keeping pace with massive social media posts is impossible. As Brandolini’s law indicates[21], correcting misinformation is laborious—often requiring one to search for related and trusted articles and write justifications—whereas creating and spreading misinformation is easy. As a consequence, a 88% of suspicious content on X did not receive any response, and 93% did not receive a high-quality response within the first hour (according to X/Twitter Community Notes as of February 2023; Supplementary Fig. S1). Even high-quality responses suffer from limited effectiveness, when they are created after rather than before initial bursts of attention to misinformation content[4, 14]. The absence of sufficient coverage also markedly diminishes the impact of correction and, as the implied truth effect suggests, may even increase the perceived accuracy of misinformation that escapes correction[22].
@@
While generative AI models such as LLMs (e.g., OpenAI’s GPT series) raise concerns that they facilitate creating misinformation, they also potentially make scaling up and accelerating misinformation’s correction possible. Recent LLMs have exhibited proficiency in generating fluent and coherent text, laying a foundation for producing explanations that the public can understand. Indeed, LLMs have revolutionized the field of AI and presented remarkable capabilities across domains and tasks[23, 24, 25]. However, accurate and trustworthy misinformation correction is inseparable from accessing up-to-date and reliable information, providing accurate references to back up claims, and addressing textual and visual information, all areas where existing LLMs fall short[26]. For instance, GPT-4[27] (as of March 2023) and MisinfoCorrect[28] lack access to timely updated knowledge and are thereby ill-equipped to combat misinformation on emerging topics. They either do not provide or “hallucinate” references, which can be fabricated or irrelevant[26, 29]. A growing body of literature has focused on retrieval-augmented LLMs, which can retrieve up-to-date information from Wikipedia or the entire Internet[30, 31, 32, 33, 34, 35, 36, 37]. However, their retrieval does not explicitly consider the factuality and bias of retrieved sources, posing risks of generating falsehood and backfiring (i.e., reinforcing rather than reducing false beliefs)[31, 38, 39]. These LLMs, as well as MisinfoCorrect[28], also struggle with counteracting multimodal misinformation due to their nonacceptance of visual inputs. Finally, understanding the quality of a correction necessitates a comprehensive evaluation of fact-checking experts due to the complexity of real-world misinformation[26]. However, prior evaluation of LLMs was conducted by non-experts on limited aspects of their generated responses (e.g., the general quality of responses[28])[37], which is insufficient to comprehensively understand their performance in identifying and explaining (in)accuracies, generated text, and references.
@@
In this article, we propose Muse, a scalable approach for multimodal misinformation correction. Muse makes use of an LLM and augments it with the ability to handle images, access timely and credible knowledge on the web, retrieve evidence that refutes or contextualizes the given content that may or may not be misinformation, and generate clear explanations with accurate and trustworthy references. Fact-checking experts comprehensively evaluate Muse-generated responses to real social media posts that potentially are misinformation, and compare them to baselines including GPT-4 and high-quality responses based on the collective efforts of laypeople. Our assessment measures the overall quality of a response, specifically defined as the explicitness, accuracy, comprehensiveness, and informativeness when it identifies and explains (in)accuracies, the relevance, factuality, fluency, coherence, and toxicity of generated text, and the reachability, relevance, and credibility of references. We find that Muse outperforms GPT-4 by 37% and even high-quality responses from laypeople by 29% in effectively and promptly responding to potential misinformation. Results demonstrate Muse’s advance when the content is textual with or without images across a broad range of domains, including politics and international affairs, economy and business, crime and law, social issues and human rights, and health and medicine.
@@
Muse is designed to automatically respond to content that potentially is misinformation. In other words, the content might be inaccurate, partially accurate, or even factually accurate but misleading, all of which are misinformation, or fully accurate as non-misinformation. The content can contain text with or without visual information. The response should identify what part(s) of the content is (in)accurate, explain why that part of the content is (in)accurate, and provide links as references. The pipeline of Muse is illustrated in Fig. 1.
@@
We start by introducing the details of Muse with a piece of text-only potential misinformation as the input. First, Muse generates queries based on an LLM from the potential misinformation (Fig. 1b; Methods). Each query acts as the input of a web search engine to access timely updated web content and obtain a list of web links directly relevant to the query (Fig. 1b; Methods). After scraping the content from these web links, Muse calculates their direct relevance to the potential misinformation and removes irrelevant web pages (Fig. 1b; Methods). Then, Muse determines the credibility of web pages by looking up their publishers’ factuality and bias ratings and selects pages with high factuality and minimal bias (Fig. 1c; Methods). Next, Muse leverages an LLM to extract text from each of the web pages as evidence. Such evidence can refute the potential misinformation, typically happening when it is misinformation with false claims, or provide additional context, which can demonstrate that the potential misinformation is accurate or part(s) of its claims are accurate (Fig. 1d; Methods). Finally, Muse generates a response to the potential misinformation by providing an LLM with the extracted pieces of evidence and their source web links (Fig. 1d; Methods).
@@
Note that content, especially content posted on social media, often contains extraneous information that does not need verification and is irrelevant to correction, including unverifiable opinions or emojis, such as the textual content of the false post in Fig. 1. Therefore, generating queries instead of simply using the post content improves web searches as a way of denoising the post content; see example queries in Fig. 1b. Meanwhile, generating multiple queries helps decompose a post, which may have multiple claims that each needs verification or correction, whereas generating one query may overlook some of the claims and hence lead to not comprehensive identifications and explanations of (in)accuracies (Supplementary Fig. S2). Another concern may arise from filtering retrieved web pages based on how relevant the page content is to the potential misinformation: theoretically, including all retrieved web pages increases the amount of extracted evidence, which may not hurt and perhaps even benefit correction. However, the increase in selected web pages drastically elevates the expense of Muse (Methods). We also observe that retrieved web pages with relatively low relevance can increase the prevalence of hallucinations when generating responses (Supplementary Fig. S3). Moreover, as illustrated in Fig. 1c, Muse filters and ranks the selected web pages by their publishers’ factuality and bias. It starts extracting evidence from pages with the highest factuality and least bias and then continues down the ranking, stopping when it has obtained sufficient refutations (i.e., at least two web pages were found to refute the misinformation) or gone through all the credible pages.
@@
When the input is a piece of multimodal (textual and visual) content, Muse first generates textual description of each image (Fig. 1a) so that the content can be handled by any LLM on downstream tasks, including query generation (Fig. 1b), evidence extraction (Fig. 1d), and response generation (Fig. 1d). Specifically, Muse augments image captioning models developed to describe an image in natural language with recognizing celebrities and optical characters based on an LLM (Fig. 1a; Methods). Compared to existing image captioning models that capture global features of images, Muse produces more informative descriptions with additional features, including identification of celebrities and embedded text—information crucial for making accurate verification and corrections[40]. For example, even a state-of-the-art image captioning model[41] may describe the visual misinformation in Fig. 1 as simply “list of banned books in Florida.” The description overlooks the listed titles of books that are essential for the visual content’s verification and correction (see more examples in Supplementary Fig. S4). In addition, Muse conducts multimodal search on the web and computes multimodal relevance to filter out irrelevant web pages (Fig. 1b; Methods).
@@
Muse’s evaluation was based on X Community Notes data. Community Notes empowers people on X, often laypeople, to collaboratively fact-check tweets, which has been shown to reduce the spread of misinformation[17]. Every laypeople’s free-response fact-check is associated with a helpfulness score by aggregating the assessments of people with diverse backgrounds, e.g., different political ideologies. A response with a sufficiently high helpfulness score is then displayed on the corresponding tweet and publicly visible[17] (Supplementary Fig. S6). We included the tweets from Community Notes with at least two responses (#=247); one has a high helpfulness score and the other has an average helpfulness score (as of February 2023; Methods). Though we do not presuppose the accuracy of the tweets in Muse’s design and evaluation, we found that more than half of the tweets are not fully (in)accurate or misleading but frequently presented in a way that combines accurate claims and inaccurate or misleading claims (Methods). We further generated responses to these tweets based on Muse and GPT-4 (as of June 2023) (Discussion; Methods). Experts in fact-checking and journalism evaluated the quality of responses by various approaches to the same tweet (Methods); they were blinded to which approach had generated each response. The evaluation contains 13 specific criteria, covering how well a response identifies and explains (in)accuracies, the quality of generated text, and the quality of references (Methods). It also contains the overall quality of a response by taking all 13 evaluation criteria into account (Methods).
@@
Results also demonstrate that Muse outperforms GPT-4 and laypeople who produce even high-helpfulness responses in the quality of generated text (Fig. 2g-k). Muse, when it augments GPT-4 with the capabilities of accessing timely updated knowledge and addressing visuals (Methods), exhibits enhanced relevance ($p=10^{-15}$; $N=460$) and factuality ($p=2\times 10^{-20}$; $N=459$) of text compared to GPT-4 without sacrificing fluency ($p=0.6$; $N=464$), coherence ($p=0.1$; $N=459$), and toxicity ($p=0.8$; $N=464$). Meanwhile, Muse-generated text is more relevant to the responded tweet ($p=2\times 10^{-30}$; $N=464$), factual ($p=4\times 10^{-6}$; $N=463$), fluent ($p=2\times 10^{-10}$; $N=464$), and coherent ($p=10^{-5}$; $N=451$) than the text of high-helpfulness responses by laypeople and additionally less toxic than the text of average-helpfulness responses by laypeople ($p=4\times 10^{-12}$; $N=462$). In particular, Muse-generated text has an average relevance score of 8.7, 18% higher than GPT-4-generated text, 21% higher than the text of high-helpfulness responses by laypeople, and 43% higher than the text of average-helpfulness responses by laypeople (Fig. 2g). 74% of Muse-generated text is completely factual, vs 59% for the text of even high-helpfulness responses by laypeople and 45% for GPT-4-generated text (Fig. 2h). Almost all of Muse-generated text does not have any mistake in the use of English (Fig. 2i) and is not biased, impolite, and provoking (Fig. 2k), and 91% is highly coherent and logical (vs 76% and 61% for laypeople, Fig. 2j).
@@
Furthermore, we observed that the quality of Muse-generated responses to textual and multimodal (textual and visual) content that potentially is misinformation is higher than responses by GPT-4 and even high-helpfulness responses by laypeople (Fig. 3a; Supplementary Fig. S7-S8). The quality of Muse-generated responses to text-based tweets is 29% higher than GPT-4-generated responses ($p=6\times 10^{-28}$; $N=310$), 33% higher than laypeople’s high-helpfulness responses ($p=10^{-44}$; $N=310$), and 65% higher than laypeople’s average-helpfulness responses ($p=10^{-67}$; $N=310$). The quality of Muse-generated responses to multimodal tweets is generally 21% higher than laypeople’s high-helpfulness responses ($p=3\times 10^{-8}$; $N=154$), 39% higher than laypeople’s average-helpfulness responses ($p=10^{-16}$; $N=154$), and 56% higher than GPT-4 ($p=6\times 10^{-18}$; $N=152$).
Finally, we found that the quality of Muse-generated responses to potential misinformation across domains including politics and international affairs, economy and business, crime and law, social issues and human rights, and health and medicine is higher than responses by GPT-4 and even high-helpfulness responses by laypeople (Methods; Fig. 3b; Supplementary Fig. S9-S13). The quality of Muse-generated responses is 35% (political and international affair; $p=10^{-16}$; $N=160$), 47% (economy and business; $p=4\times 10^{-11}$; $N=76$), 44% (crime and law; $p=3\times 10^{-8}$; $N=76$), 31% (social issues and human rights; $p=0.0002$; $N=60$), and 27% (health and medicine; $p=3\times 10^{-5}$; $N=48$) higher than GPT-4-generated responses. The quality of Muse-generated responses is 33% (political and international affair; $p=10^{-21}$; $N=160$), 27% (economy and business; $p=10^{-8}$; $N=76$), 28% (crime and law; $p=3\times 10^{-9}$; $N=76$), 38% (social issues and human rights; $p=7\times 10^{-10}$; $N=60$), and 25% (health and medicine; $p=4\times 10^{-5}$; $N=48$) higher than even high-helpfulness responses by laypeople.

### v4

@@ Abstract
Real-world misinformation, often multimodal, can be partially correct and factual but misleading by cherry-picking, conflating correlation with causation, and using other tactics. Such misinformation is severely understudied, challenging to address, and harms many social domains, including democracy, economics, justice, human rights, and public health, particularly on social media, where it can spread rapidly. High-quality and timely correction of misinformation that identifies and explains its inaccuracies and accuracies has been shown to effectively reduce false beliefs. Despite the wide acceptance of manual correction, it is difficult to be timely and scalable, a concern as technologies like large language models (LLMs) make misinformation easier to produce. LLMs also have versatile capabilities that could accelerate misinformation correction—however, they struggle due to a lack of recent information, a tendency to produce false content, and limitations in addressing multimodal information. We propose Muse, an LLM augmented with access to and credibility evaluation of up-to-date information. By retrieving evidence as refutation or supporting context, Muse identifies and explains (in)accuracies in a piece of content—not presupposed to be misinformation—with references. It conducts multimodal retrieval and interprets visual content to verify and correct multimodal content. Given the absence of a systematic and comprehensive evaluation approach, we propose and define 13 dimensions of misinformation correction quality, ranging from the accuracy of identifications and factuality of explanations to the relevance and credibility of references. Then, fact-checking experts correspondingly evaluate responses to social media content that are not presupposed to be (non-)misinformation but broadly include incorrect, partially correct, and correct posts that may or may not be misleading. The results demonstrate Muse’s ability to write high-quality responses to potential misinformation—across modalities, tactics, domains, political leanings, and for information that has not previously been fact-checked online—within minutes of its appearance on social media. Overall, Muse outperforms GPT-4 by 37% and even high-quality responses from laypeople by 29%. Our work provides a general methodological and evaluative framework to correct misinformation at scale. It reveals LLMs’ potential to help combat real-world misinformation effectively and efficiently.
@@ Introduction
Misinformation has far-reaching and detrimental effects on individuals and society[1, 2, 3, 4, 5]. It erodes public trust in government, decreases civil engagement in elections, and has been viewed as a threat to democracy[6, 3, 7]. Evidence has demonstrated that election misinformation helped fuel the January 6th attack on the U.S. Capitol, where five people died and more than 100 police officers were injured[8]. Misinformation also drastically increases during outbreaks and disasters, as seen with the “infodemic” during the COVID-19 pandemic, which significantly increased vaccine hesitancy[9, 10]. The result is tragic, where COVID-19 vaccines could have prevented at least 232,000 COVID-19-associated deaths between May 2021 and September 2022 in the U.S. alone[11]. Concerns over misinformation on social media have been particularly significant[12, 13, 14], as the social media context interferes with truth discernment, where users post content without professional moderation and often consume news in a hasty and distracted way[15, 16].
@@
Fortunately, high-quality and timely correction of misinformation, which identifies what part(s) of its content is or is not accurate and explains why that part of the content is (in)accurate with references, has been shown to effectively reduce the spread of misinformation and false beliefs[17, 18, 4, 19]. While domain experts (e.g., on FactCheck.org) and groups of laypeople (e.g., on X Community Notes, formerly Twitter Birdwatch) have played pivotal roles in correcting misinformation[19, 20, 17], keeping pace with massive social media posts is impossible. As Brandolini’s law indicates[21], correcting misinformation is laborious—often requiring domain knowledge, information and media literacy, and explanatory skills—whereas creating and spreading misinformation is easy. As a consequence, 88% of suspicious content on X did not receive any response, and 93% did not receive a high-quality response within the first hour (according to X/Twitter Community Notes as of February 2023; Supp. Fig. S1). Even high-quality responses suffer from limited effectiveness when they are created after rather than before initial bursts of attention to misinformation content[4, 14]. The absence of sufficient coverage also markedly diminishes the impact of correction and, as the implied truth effect suggests, may even increase the perceived accuracy of misinformation that escapes correction[22].
@@
While generative AI models like LLMs raise concerns that they facilitate creating misinformation, they also potentially make scaling up and accelerating misinformation’s correction possible. Recent LLMs have exhibited proficiency in generating fluent and coherent text, laying a foundation for producing explanations that the public can understand. Indeed, LLMs have revolutionized the field of AI and presented remarkable capabilities across domains and tasks[23, 24, 25]. However, accurate and trustworthy misinformation correction is inseparable from accessing up-to-date and reliable information, providing accurate references to back up claims, and addressing textual and visual information, all areas where existing LLMs fall short[26]. Specifically, GPT-4[27] (as of March 2023), Llama-3 (as of July 2024)[28], MisinfoCorrect[29], MADR[30], JustiLM[31], and many other LLMs lack access to constantly growing and changing knowledge. These models are thereby ill-equipped to combat misinformation on emerging topics. They either do not provide or “hallucinate” references, which can be fabricated or irrelevant[26, 32]. A growing body of literature has focused on retrieval-augmented LLMs, which can retrieve up-to-date information from Wikipedia or the entire Internet[33, 34, 35, 36, 37, 38, 39]. However, their retrieval does not explicitly consider the factuality and bias of retrieved sources, posing risks of generating falsehood and backfiring (i.e., reinforcing rather than reducing false beliefs)[40, 41, 42]. Every above LLM also struggles with counteracting multimodal misinformation due to their nonacceptance of visual inputs.
Real-world misinformation commonly encompasses content that is partially correct and even factual but misleading through various tactics and across domains[2, 3]. Such misinformation can be more difficult to recognize, prevalent, and harmful[2, 3]. Existing LLM and other AI models to tackle misinformation focus narrowly on it that is factually incorrect (e.g., MADR[30], FOLK[38], and ProgramFC[43]), within a single domain (e.g., politics[30] or COVID-19[29, 39]), or employing one specific tactic (e.g., repurposing authentic images with fabricated stories[44]). A general AI framework to practically identify and correct misinformation is demanded but absent. Evaluating the quality of a correction presented in natural language is another core challenge. Related studies heavily rely on automatic evaluation, which calculates the similarity between model-generated and “ground-truth” text using metrics such as ROUGE[31, 44]. However, what qualifies a correction as the ground truth is unclear[29], and existing metrics struggle to capture content factuality, especially when compared to fact-checking experts[26]. The need for an accurate and comprehensive understanding of a model’s performance in identifying and explaining (in)accuracies, generated text, and references is recognized, but a solution is still lacking.
@@
In this article, we propose Muse, a scalable approach for multimodal misinformation correction. Muse makes use of an LLM and augments it with the ability to handle images, access timely and credible knowledge on the web, retrieve evidence that refutes or contextualizes the given content that may or may not be misinformation, and generate clear explanations with accurate and trustworthy references. Muse is a nonparametric approach and therefore easily and cheaply updated, especially compared to parametric language models. Fact-checking experts comprehensively evaluate Muse-generated responses to real social media posts that potentially are misinformation, and compare them to baselines including GPT-4 and high-quality responses based on the collective efforts of laypeople. Our assessment measures the overall quality of a response, specifically defined as the explicitness, accuracy, comprehensiveness, and informativeness when it identifies and explains (in)accuracies; the relevance, factuality, fluency, coherence, and toxicity of generated text; and the reachability, relevance, and credibility of references. We find that Muse outperforms GPT-4 by 37% and even high-quality responses from laypeople by 29% in accurately and promptly responding to potential misinformation. Results demonstrate Muse’s advance no matter whether the potential misinformation content is textual with or without images; fabricates a story, misinterprets or misrepresents facts, lacks context, implies false or oversimplified causation, uses loaded language, presents false or biased data, has improper analogy or equivalence, and adopts other tactics; across a broad range of domains, including politics and international affairs, economy and business, crime and law, social issues and human rights, and health and medicine; leans liberal or conservative; or whether the content has previously fact-checked online or not.
@@
Muse is designed to automatically respond to content that potentially is misinformation. In other words, the content might be (partially) inaccurate or factually accurate but misleading, all of which are misinformation, or fully accurate as non-misinformation. The content can contain text with or without visuals. The response should identify what part(s) of the content is or is not accurate, explain why that part of the content is or is not accurate, and provide links as references. Overall, Muse features three modules to reliably and efficiently address generalized misinformation: (1) a highly parallelizable relevance- and credibility-aware evidence retriever, (2) an LLM-based evidence-assisted response generator, and (3) an image describer that enables text-only LLMs to jointly interpret multimodal signals from the input.
@@
Fig. 1 illustrates Muse’s pipeline. We start by introducing the details of Muse with a piece of text-only potential misinformation as the input. First, Muse generates queries based on an LLM from the potential misinformation (Fig. 1b; Methods). Each query acts as the input of a web search engine to access timely updated web content and obtain a list of web links directly relevant to the query (Fig. 1b; Methods). After scraping the content from these web links, Muse calculates their direct relevance to the potential misinformation and removes irrelevant web pages (Fig. 1b; Methods). Then, Muse determines the credibility of web pages by looking up their publishers’ factuality and bias ratings and selects pages with high factuality and minimal bias (Fig. 1c; Methods). Next, Muse leverages an LLM to extract text from each of the web pages as evidence. Such evidence can refute the potential misinformation, typically happening when it is misinformation with false claims, or provide additional context, which can demonstrate that the potential misinformation is accurate or part(s) of its claims are accurate (Fig. 1d; Methods). Finally, Muse generates a response to the potential misinformation by providing an LLM with the extracted pieces of evidence and their source web links (Fig. 1d; Methods).
@@
Note that content, especially content posted on social media, often contains extraneous information that does not need verification and is irrelevant to correction, including unverifiable opinions or emojis, such as the textual content of the false post in Fig. 1. Therefore, generating queries instead of simply using the post content improves web searches as a way of denoising the post content; see example queries in Fig. 1b. Meanwhile, generating multiple queries helps decompose a post, which may have multiple claims that each needs verification or correction, whereas generating one query may overlook some of the claims and hence lead to not comprehensive identifications and explanations of (in)accuracies (Supp. Fig. S2). Another concern may arise from filtering retrieved web pages based on how relevant the page content is to the potential misinformation: theoretically, including all retrieved web pages increases the amount of extracted evidence, which may not hurt and perhaps even benefit correction. However, the increase in selected web pages drastically elevates the expense of Muse (Methods). We also observe that retrieved web pages with relatively low relevance can increase the prevalence of hallucinations when generating responses (Supp. Fig. S3). Moreover, as illustrated in Fig. 1c, Muse filters and ranks the selected web pages by their publishers’ factuality and bias. It starts extracting evidence from pages with the highest factuality and least bias and then continues down the ranking, stopping when it has obtained sufficient refutations (i.e., at least two web pages were found to refute the misinformation) or gone through all the credible pages.
@@
When the input is a piece of multimodal (textual and visual) content, Muse first generates textual description of each image (Fig. 1a) so that the content can be handled by any LLM on downstream tasks, including query generation (Fig. 1b), evidence extraction (Fig. 1d), and response generation (Fig. 1d). Specifically, Muse augments image captioning models developed to describe an image in natural language with recognizing celebrities and optical characters based on an LLM (Fig. 1a; Methods). Compared to existing image captioning models that capture global features of images, Muse produces more informative descriptions with features crucial for making accurate verification and corrections[45] (Methods). For example, even a state-of-the-art image captioning model[46] may describe the visual misinformation in Fig. 1 as simply “list of banned books in Florida.” The description overlooks the listed titles of books that are essential for the visual content’s verification and correction (see more examples in Supp. Fig. S4). In addition, Muse conducts multimodal search on the web and computes multimodal relevance to filter out irrelevant web pages (Fig. 1b; Methods).
@@
Muse’s evaluation was based on X Community Notes data. Community Notes empowers people on X, often laypeople, to collaboratively fact-check tweets, which has been shown to reduce the spread of misinformation[17]. Every laypeople’s free-response fact-check is associated with a helpfulness score by aggregating the assessments of people with diverse backgrounds, e.g., different political ideologies. A response with a sufficiently high helpfulness score is then displayed on the corresponding tweet and publicly visible[17] (Supp. Fig. S6). We included the tweets from Community Notes with at least two responses ($n=247$, posted from February 2021 to February 2023); one has a high helpfulness score and the other has an average helpfulness score (as of February 2023; Methods). Though we do not presuppose the accuracy of the tweets in Muse’s design and evaluation, we found that more than half of the tweets are not fully (in)accurate or misleading but frequently presented in a way that combines accurate claims and inaccurate or misleading claims (Methods). We further generated responses to these tweets based on Muse and GPT-4 (as of June 2023) (Discussion; Methods). Experts in fact-checking and journalism evaluated the quality of responses by various approaches to the same tweet (Methods); they were blinded to which approach had generated each response. The evaluation contains 13 specific criteria, covering how well a response identifies and explains (in)accuracies, the quality of generated text, and the quality of references (Methods). It also contains the overall quality of a response by taking all 13 evaluation criteria into account (Methods).
@@
Results also demonstrate that Muse outperforms GPT-4 and laypeople who produce even high-helpfulness responses in the quality of generated text (Fig. 2g-k). Muse, when it augments GPT-4 with the capabilities of accessing timely updated knowledge and addressing visuals (Methods), exhibits enhanced relevance ($p=10^{-15}$; $N=460$) and factuality ($p=2\times 10^{-20}$; $N=459$) of text compared to GPT-4 without sacrificing fluency ($p=0.6$; $N=464$), coherence ($p=0.1$; $N=459$), and toxicity ($p=0.8$; $N=464$). Meanwhile, Muse-generated text is more relevant to the responded tweet ($p=2\times 10^{-30}$; $N=464$), factual ($p=4\times 10^{-6}$; $N=463$), fluent ($p=2\times 10^{-10}$; $N=464$), and coherent ($p=10^{-5}$; $N=451$) than the text of high-helpfulness responses by laypeople and additionally less toxic than the text of average-helpfulness responses by laypeople ($p=4\times 10^{-12}$; $N=462$). Muse-generated text has an average relevance score of 8.7, 18% higher than GPT-4-generated text, 21% higher than the text of high-helpfulness responses by laypeople, and 43% higher than the text of average-helpfulness responses by laypeople (Fig. 2g). 74% of Muse-generated text is completely factual, vs 59% for the text of even high-helpfulness responses by laypeople and 45% for GPT-4-generated text (Fig. 2h). Almost all of Muse-generated text does not have any mistake in the use of English (Fig. 2i) and is not biased, impolite, and provoking (Fig. 2k), and 91% is highly coherent and logical (vs 76% and 61% for laypeople, Fig. 2j).
@@
We further analyzed the robustness of Muse in addressing diverse forms of real-world misinformation, which may vary in modality, difficulty to correct, political leaning, domain, and adopted tactics. We first observed that the quality of Muse-generated responses to textual and multimodal (textual and visual) content is consistently higher than responses by GPT-4 and even high-helpfulness responses by laypeople, with a margin of 21% or higher (Fig. 3a). GPT-4 can be comparable to laypeople who write high-helpfulness responses when responding to text-only tweets. However, its performance significantly declines when responding to multimodal tweets, often resulting in a quality lower than laypeople who produce average-helpfulness responses.
Second, we removed tweets where Muse’s response references at least one related fact-checking article, accounting for 16% of all tweets. Notably, in only 6% of all tweets, Muse’s response exclusively references related fact-checking article(s). The quality of Muse-generated responses to these tweets can be more associated with Muse’s ability to summarize retrieved professional fact-checking articles, rather than fact-checking such content from scratch. We found that the quality of Muse-generated responses to content that has not been fact-checked online is at least 28% higher than responses by GPT-4 and even high-helpfulness responses by laypeople (Fig. 3b). The average quality score of Muse-generated responses for these potentially more challenging tweets is 8.0, close to its 8.1 average score for all tweets.
Third, LLMs have exhibited political bias [47], which may diminish their ability to debunk misinformation that aligns with their political stance. We annotated the political leaning of tweets and found that GPT-4, which has an observed liberal bias [47], indeed performs 5% worse in correcting misinformation that aligns with liberal ideology compared to conservative ideology (Fig. 3c). Nevertheless, the quality of Muse-generated responses to both liberal and conservative content has the same average score of 8.3. This score is consistently higher than GPT-4-generated responses and even high-helpfulness responses from laypeople by at least 26%.
Finally, we annotated the domain of tweets and tactics used to make them or part of them false or misleading. Results in Fig. 3d-e indicate that the quality of Muse-generated responses to potential misinformation across different domains and tactics maintains an average score around 8, consistently higher than responses by GPT-4 and even high-helpfulness responses from laypeople, with a margin of at least 19%. Here, domains broadly include politics and international affairs, economy and business, crime and law, social issues and human rights, and health and medicine. Tactics primarily include fabricating, e.g., a story with or without an authentic image, or using a digitally fabricated or altered screenshot or photo; misinterpreting or misrepresenting someone’s claim, a symbol, a policy, and others; lacking context; implying false or oversimplified causation, such as conflating correlation with causation; using loaded language; presenting false, partial, or biased data, such as cherry-picking; and having improper analogy or equivalence (see examples in Supp. Fig. S7).